OpenStudy (anonymous):

Given the data points (1, 4), (2, 9), and (4, 19), which of the following equations describes the best-fit line? A: y = -1 + 5x B: y = 5 +1x C: y = 2 + 5x D: y = 1 + 5x

OpenStudy (anonymous):

(x1,y1) = (1,4) (x2,y2) = (2,9) (x3,y3) = (4,19) slope of the line,m = y2-y1/x2-x1 = y3-y2/x3-x2 point slope form = (y-y1) = m(x-x1) then make above eq into y = mx+c